OpenStudy (kels2153):

cos^2(x)=1/2 My main question is how to get cos^2 to cos?

OpenStudy (kels2153):

I know that \[\cos^2=1-\sin^2\] but I am stuck after that.

OpenStudy (amedel):

It doesn't look like you need to use an identity here. Whatever cos(x) is, squared, is equal to 1/2. So try to get rid of the exponent.

OpenStudy (amedel):

If it helps, substitute a = cos(x). Then we have: a^2 = 1/2

OpenStudy (kels2153):

My mind has been enlightened. Thank you so much!

OpenStudy (amedel):

No problem, glad to help!